What's More



Let's see how well you understood our discussion. At this point, I want you to



solve the following problems. Show your complete solution by following the step-by-



step procedure.



1. The average number of milligrams (mg) of cholesterol in a cup of a certain brand



of ice cream is 660 mg, the standard deviation is 35 mg. Assume the variable is



normally distributed.



a. If a cup of ice cream is selected, what is the probability that the cholesterol



content will be more than 670 mg?



b. If a sample of 10 cups of ice cream is selected, what is the probability that



the mean of the sample will be larger than 670 mg?



2. In a study of the life expectancy of 400 people in a certain geographic region, the



mean age at death was 70 years, and the standard deviation was 5.1 years. If a



sample of 50 people from this region is selected, what is the probability that the



mean life expectancy will be less than 68 years?




1
Expert's answer
2022-04-10T15:32:34-0400

1.

a. "(X>670)=1\u2212P(X\u2264670)\n=1-P(Z\\leq\\dfrac{670-660}{35})=1\u2212P(Z\u2264 \\frac{\n \n670\u2212660}{35}\n\u200b\n )\n\\approx1-P(Z\\leq0.2857)=1-0.61409=0.38591"

b."(X>670)=1\u2212P(X\u2264670)\n=1-P(Z\\leq\\dfrac{670-660}{35\/\\sqrt{10}})=1\u2212P(Z\u2264 \\frac{\n \n670\u2212660}{35\/\\sqrt{10}}\n\u200b\n )\n\\approx1-P(Z\\leq0.9035)\\approx0.1831"

2.

"P(X<68)=P(Z<\frac{68-70}{5.1/\sqrt{50}})=P(Z<-2.77)=0.0028"